On the body with mass m = 1kg, moving in a horizontal plane, the force F = 10N acts, parallel to the horizontal plane. The body moves with acceleration a = 6 m / s². The coefficient of friction of the body with the plane is: 

 A) 0.1 

 B) 0.2 

 D) 0.6 

 C) 0.4


1
Expert's answer
2021-03-14T19:19:21-0400

Let's apply the Newton's Second Law of Motion:


"F-F_{fr}=ma,""F-\mu_k mg=ma,""\mu_k=\dfrac{F-ma}{mg},""\mu_k=\dfrac{10\ N-1\ kg\cdot6\ \dfrac{m}{s^2}}{1\ kg\cdot9.8\ \dfrac{m}{s^2}}=0.4"